# The Art of Imitation: Learning Long-Horizon Manipulation Tasks from Few Demonstrations

### v1

@@ I Introduction
We propose Task Auto-Parameterized And Skill Segmented GMM TAPAS-GMM, illustrated in Fig. 1. Our method addresses this set of challenges synergistically. First, we propose to refactorize the robot’s velocity into direction and magnitude to better capture the underlying data distribution. We then model both using Riemannian GMMs [10] which are not constrained by the Euclidean assumptions. Second, we segment the task demonstrations into skills using the velocity information. The segmentation both aligns the trajectories temporally and enables learning skill models with shorter trajectories and lower data dimensionality. Additionally, the temporal alignment enables us to use time as a powerful inductive bias for the initialization of our model. If the trajectories are aligned, then simple binning of the trajectories over the time dimension yields an excellent initialization. Third, we use a pretrained visual encoder network [12] to generate object keypoints that generalize across object instances and task environments [4]. This way, we generate a set of candidate task parameters from RGB-D observations. Finally, for selecting the task-relevant parameters from these candidates, we offer an approach that scales well to large numbers of candidates. Our proposed solutions have synergistic effects. For example, combining our parameter selection and skill segmentation further reduces the data dimensionality, as a given task parameter is often only relevant for some skill and not the whole long-horizon task.
We evaluate our method on a diverse set of manipulation tasks, both in RLBench [13] and on a real robot, including articulated objects, as well as high-precision, and long-horizon tasks. We experimentally disentangle policy learning and representation learning, as well as ablating the major components of our approach. TAPAS-GMM learns long-horizon manipulation policies from as little as five demonstrations. We show generalization across object positions, object instances, visual clutter, and task environments. Finally, we demonstrate how the learned skills can be recombined in novel ways.
@@
    rigorous evaluation of our approach’s ability to overcome the problems outlined above, and its efficacy for policy learning, both in simulation and on a real robot.
@@ II Related Work
Parameterized Motions: GMMs [5] learn robotic manipulation tasks from only a handful of demonstrations. Popular alternatives are Dynamic Motion Primitives (DMP) [14] and Probabilistic Motion Primitives (ProMP) [15]. However, all three methods require additional techniques to adapt to novel task instances. Calinon et al. [6] identify three classes of adaptation methods: model database, parametric, and multi-stream, to which we amend two-stage and point-based approaches. Model database approaches [16] learn one model per demonstration, that they retrieve and combine based on the associated parameters. In contrast, parametric models [17] learn a single, parameterized model. Two-stage approaches first learn a model encoding a primitive motion, that they adapt in a second stage, e.g. using reinforcement learning [18] or Wasserstein gradients [19]. However, this requires a second online learning phase. Whereas, multi-stream approaches [20] model the same set of demonstrations from the perspective of multiple coordinate frames, called task-parameters. They generalize via the transformation and combination of these local models. Task-parameterized Gaussian Mixture Models (TP-GMM) [6, 10, 21] are multi-stream methods, leveraging a single expectation-maximization procedure. Finally, point-based approaches [9, 22] warp the underlying model to explicit endpoints and waypoints.
@@
Sample-Efficient Long-Horizon Manipulation: GMMs are typically used to model full trajectories [18], with some work focusing on solving long-horizon tasks by sequencing learned GMMs in a bottom-up manner [21]. In contrast, we focus on directly learning long-horizon tasks by segmenting and sequencing skills top-down from complex task demonstrations. Some deep learning based approaches [25, 26] simplify long-horizon problems, by only predicting a set of key poses and using motion planning in between. However, by throwing away the trajectory information between key poses, they cannot produce constraint motions such as opening a cabinet. Moreover, they can only predict key poses in the demonstrated order, whereas our approach further produces a set of reusable and re-combinable skills.
@@ III Background
In this section, we summarize foundational work on Gaussian Mixture Models for robotic manipulation, including Hidden Markov Models (HMM), task-parameterization, and Riemannian GMMs. We provide additional details including full equations in Sec. S.1 of the supplementary material.
[…]
@@ III-A Gaussian Mixture Model
A Gaussian Mixture Model (GMM) with $K$ Gaussian components is defined as $\mathcal{G}=\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$. $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\Sigma}_{k}$ denote the $k$-th component’s mean and covariance matrix and $\pi_{k}$ denotes its prior probability. In the context of Learning from Demonstrations, GMMs are employed to estimate the joint distribution of a disjoint set of input variables $\mathcal{I}$ and output variables $\mathcal{O}$, from a set of task demonstrations $\mathcal{D}$ as
@@
III-C Hidden Markov Model
The GMM’s posterior $p_{\mathcal{G}}(\boldsymbol{o}\mid\boldsymbol{i})$ in Eq. (2) only depends on the priors $\pi_{k}^{i}$ and the conditioned Gaussians, not on any previous observations. The Hidden Markov Model (HMM) [29] augments the GMM by associating each Gaussian with a state $k$ and modeling the transition probability $a_{kl}$ from state $k$ to state $l$. I.e. =$\mathcal{H}=\{\pi_{k},\{a_{kl}\}_{l=1}^{K},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$. This helps the model to differentiate between components that are close together in the input space, yet model different parts of a task or movement.. For a sequence of observations $\{\boldsymbol{i}_{t}\}_{t=0}^{T}$, the computation of the prior thus changes to
$\pi_{k}^{i_{t}}=\frac{\left(\sum_{l=1}^{K}\pi_{l}^{i_{t-1}}a_{lk}\right)\mathcal{N}(\boldsymbol{i}_{t}\mid\boldsymbol{\mu}_{k}^{\mathcal{I}},\boldsymbol{\Sigma}_{k}^{\mathcal{I}})}{\sum_{m=1}^{K}\left[\left(\sum_{l=1}^{K}\pi_{l}^{i_{t-1}}a_{lm}\right)\mathcal{N}(\boldsymbol{i}_{t}\mid\boldsymbol{\mu}_{m}^{\mathcal{I}},\boldsymbol{\Sigma}_{m}^{\mathcal{I}})\right]}.$ (3)
For clarity, we continue referring to our models as GMMs, even though all of them use the HMM extension.
III-D Task-Parameterization
Task-Parameterized GMMs [6] model the movement from the perspective of multiple coordinate frames, called task-parameters, in one joint GMM [10]. These frames are attached to task-relevant objects, but may also include the robot’s frame and a world frame. Let $\boldsymbol{x}(f)$ denote the end-effector position in the frame of task-parameter $f$. We discuss the extension to full poses $\boldsymbol{\xi}(f)$ in Sec. III-E. For the time-driven model and $F$ task-paramaters, we model =$\mathcal{I}=\{t\}\text{ and }\mathcal{O}=\{\boldsymbol{x}(f)\}_{f=1}^{F}$. In contrast, for the state-driven model, we define =$\mathcal{I}=\{\boldsymbol{x}(f)\}_{f=1}^{F}\text{ and }\mathcal{O}=\{\dot{\boldsymbol{x}}(f)\}_{f=1}^{F}$. Akin to the vanilla GMM, we fit all parameters jointly using EM. Next, we compute the per-frame marginal model for each frame $1\leq f\leq F$. For example, in the state-driven case we estimate
$p_{f}(\mathcal{I},\mathcal{O}\mid\mathcal{G})=p(\boldsymbol{x}(f),\dot{\boldsymbol{x}}(f)\mid\mathcal{G}).$ (4)
To adapt to a given set of task parameters $\Theta=\{\boldsymbol{A}_{f},\boldsymbol{b}_{f}\}_{f=1}^{F}$, we transform the per-frame marginals to the world frame. $\boldsymbol{b}_{f}$ denotes the origin of frame $f$ and $\boldsymbol{A}_{f}$ its rotation matrix. Let
$p_{f}(\mathcal{I},\mathcal{O}\mid\mathcal{G})=\left\{\pi_{k}^{f},\boldsymbol{\mu}_{k}^{f},\boldsymbol{\Sigma}_{k}^{f}\right\}_{k=1}^{K}$ (5)
be a marginal. Then its transformed marginal is given by
$\begin{split}p_{f}(\mathcal{I},\mathcal{O}\mid\mathcal{G},\Theta)&=\left\{\pi_{k}^{f},\hat{\boldsymbol{\mu}}^{f}_{k},\hat{\boldsymbol{\Sigma}}^{f}_{k}\right\}_{k=1}^{K}\\ \hat{\boldsymbol{\mu}}^{f}_{k}&=\boldsymbol{A}_{f}\boldsymbol{\mu}_{k}^{f}+\boldsymbol{b}_{f}\\ \hat{\boldsymbol{\Sigma}}^{f}_{k}&=\boldsymbol{A}_{f}\boldsymbol{\Sigma}_{k}^{f}\boldsymbol{A}_{f}^{T}.\end{split}$ (6)
@@ III-D Task-Parameterization
III-E Riemannian Manifolds
We model the robot’s end-effector pose $\boldsymbol{\xi}=\begin{bmatrix}\boldsymbol{x}&\boldsymbol{\theta}\end{bmatrix}^{T}$ on the Riemmanian Manifold $\mathcal{M}_{\text{pose}}=\mathbb{R}^{3}\times\mathcal{S}^{3}$. Whereas, GMMs typically assume Euclidean data, the naive Euclidean mean of a set of quaternions is generally not a valid quaternion. Instead of projecting the naive solution back onto the manifold, we solve the problem rigorously using Riemannian geometry [10]. Each Riemannian manifold $\mathcal{M}$ is differentiable, i.e., at each point $\mathbf{x}$, it has a tangent space $\mathcal{T}_{\mathbf{x}}\mathcal{M}$, that is a real vector space. The logarithmic map $\text{Log}_{\mathbf{x}}:\mathcal{M}\to\mathcal{T}_{\mathbf{x}}\mathcal{M}$ straightens the manifold’s curvature while preserving geodesic distance. The inverse is the exponential map $\text{Exp}_{\mathbf{x}}:\mathcal{T}_{\mathbf{x}}\mathcal{M}\to\mathcal{M}$. Both depend on the base $\mathbf{x}$, at which they are applied. Following Zeestraten [10], we formulate all previously Euclidean operations in EM and GMR as non-linear Maximum Likelihood Estimation problems. For example, to compute the center of mass $\boldsymbol{\mu}$ of a Riemannian Gaussian, we initialize the procedure with some estimate $\hat{\boldsymbol{\mu}}$ and solve the MLE using the following iterative process.
  1. 1.
    Tangent project the data using $\text{Log}_{\hat{\boldsymbol{\mu}}}$.
  2. 2.
    Update $\Delta\in\mathcal{T}_{\hat{\boldsymbol{\mu}}}\mathcal{M}$ using a Gauss-Newton step.
  3. 3.
    Back-project the update $\hat{\boldsymbol{\mu}}\leftarrow\text{Exp}_{\hat{\boldsymbol{\mu}}}(\Delta)$.
In practice, only a handful of iterations suffice for $\hat{\boldsymbol{\mu}}$ to converge. Fig. 2 illustrates the procedure. All other operations such as conditioning can be solved in a similar manner [10]. Finally, we use the concept of parallel transport to implement the affine frame transformations in Eq. (6) for the full pose. We provide additional details in Sec. S.2 of the supplementary material. Zeestraten provides an excellent introduction into TP-GMMs on Riemannian Manifolds [10]. Note, that we found parallel transport to sometimes introduce unwanted correlations into the model’s covariance, leading to faulty predictions. We discuss the problem and our proposed solution in detail in Sec. S.3 of the supplementary material.
@@ IV-A Action Factorization
In Sec. III-B, we discussed that solving manipulation tasks can require modeling the end-effector’s velocity $\dot{\boldsymbol{\xi}}$ and not only its pose $\boldsymbol{\xi}$, e.g. if time-independence is required. Fig. 3 illustrates that naively modeling $\dot{\boldsymbol{\xi}}$ on the manifold $\mathcal{M}_{\text{pose}}=\mathbb{R}^{3}\times\mathcal{S}^{3}$ (introduced in Sec. III-E) is difficult because velocities are not Euclidean. Instead, they are vector quantities defined by a direction and magnitude. We therefore propose to factorize the end-effector velocity $\dot{\boldsymbol{\xi}}$ as
@@
Our insight is three-fold. First, in the task-parameterized setting, only the direction of the velocities varies between different fixed coordinate frames, not their magnitude. Thus, we reduce the dimensionality of the problem by modeling the magnitudes globally. Second, because velocities are not Euclidean, their distributions are easier to estimate when factorized. To model the factorized velocities, we leverage the Riemannian toolbox introduced in Sec. III-E. Third, we leverage the isolated velocities to automatically segment complex task trajectories into the involved skills. We found this to drastically improve the quality and transferability of the learned model and discuss it in detail in Sec. IV-C. Fig. 3 illustrates how the velocities are easier to model in the factorized representation. We provide additional details on how to implement the frame transformations for $\mathcal{S}^{2}$ in Sec. S.2 of the supplementary material.
@@ IV-C Skill Segmentation
Robot manipulation tasks often involve a set of skills. The pick and place task in Fig. 5 consists of four skills. 1. Align the gripper with the first object. 2. Grasp it. 3. Align the first object with the second object. 4. Place the first object. All skills have distinct dynamics, so it is difficult to model them jointly in a single state-driven GMM. Furthermore, the skills have different relative durations across task instances, leading to poor temporal alignment across demonstrations. Fig. 4 illustrates that as a consequence they are difficult to model with a single time-driven GMM.
Therefore, we propose to automatically segment the full demonstration trajectories into a sequence of skill demonstrations and to learn one model per skill. This allows us to temporally align the skills and to learn local dynamics, specific to the skills. The skill model can be time-driven or state-driven. Furthermore, it allows us to reuse the learned skills by sequencing them in novel ways. Lastly, in Sec. IV-E and Sec. IV-F we discuss its synergistic effects with our initialization strategy and our automatic task-parameterization, respectively. To segment the trajectories, we threshold the action magnitudes introduced in Sec. IV-A. Observations with near-zero action magnitude yield a set of segmentation point candidates that we then cluster across the set of demonstrations using density-based clustering. Afterward, we fit one TP-HMM per skill.

### v2

@@ I Introduction
We propose Task Auto-Parameterized And Skill Segmented GMM TAPAS-GMM, illustrated in Fig. 1. Our method addresses this set of challenges synergistically. First, we propose to refactorize the robot’s velocity into direction and magnitude to better capture the underlying data distribution. We then model both using Riemannian GMMs [10], which are not constrained by the Euclidean assumptions. Second, we segment the task demonstrations into skills using the velocity information. The segmentation both aligns the trajectories temporally and enables learning skill models with shorter trajectories and lower data dimensionality. Additionally, the temporal alignment enables us to use time as a powerful inductive bias for the initialization of our model. If the trajectories are aligned, then simple binning of the trajectories over the time dimension yields an excellent initialization. Third, we use a pretrained visual encoder network [12] to generate object keypoints that generalize across object instances and task environments [4]. This way, we generate a set of candidate task parameters from RGB-D observations. Finally, for selecting the task-relevant parameters from these candidates, we offer an approach that scales well to large numbers of candidates. Our proposed solutions have synergistic effects. For example, combining our parameter selection and skill segmentation further reduces the data dimensionality, as a given task parameter is often only relevant for some skill and not the whole long-horizon task.
We evaluate our method on a diverse set of manipulation tasks, both in RLBench [13] and on a real robot, including articulated objects, as well as high precision and long-horizon tasks. We experimentally disentangle policy learning and representation learning, as well as ablating the major components of our approach. TAPAS-GMM learns long-horizon manipulation policies from as little as five demonstrations. We show generalization across object positions, object instances, visual clutter, and task environments. Finally, we demonstrate how the learned skills can be recombined in novel ways.
@@
    rigorous evaluation of our approach’s ability to overcome the problems outlined above and its efficacy for policy learning, both in simulation and on a real robot.
@@ II Related Work
Parameterized Motions: GMMs [5] learn robotic manipulation tasks from only a handful of demonstrations. Popular alternatives are Dynamic Motion Primitives (DMP) [14] and Probabilistic Motion Primitives (ProMP) [15]. However, all three methods require additional techniques to adapt to novel task instances. Calinon et al. [6] identify three classes of adaptation methods: model database, parametric, and multi-stream, to which we amend two-stage and point-based approaches. Model database approaches [16] learn one model per demonstration, that they retrieve and combine based on the associated parameters. In contrast, parametric models [17] learn a single, parameterized model. Two-stage approaches first learn a model encoding a primitive motion that they adapt in a second stage, e.g.,  using reinforcement learning [18] or Wasserstein gradients [19]. However, this requires a second online learning phase. Whereas, multi-stream approaches [20] model the same set of demonstrations from the perspective of multiple coordinate frames, called task-parameters. They generalize via the transformation and combination of these local models. Task-parameterized Gaussian Mixture Models (TP-GMM) [6, 10, 21] are multi-stream methods, leveraging a single expectation-maximization procedure. Finally, point-based approaches [9, 22] warp the underlying model to explicit endpoints and waypoints.
@@
Sample-Efficient Long-Horizon Manipulation: GMMs are typically used to model full trajectories [18], with some work focusing on solving long-horizon tasks by sequencing learned GMMs in a bottom-up manner [21]. In contrast, we focus on directly learning long-horizon tasks by segmenting and sequencing skills top-down from complex task demonstrations. Some deep learning based approaches [25, 26] simplify long-horizon problems by only predicting a set of key poses and using motion planning in between. However, by throwing away the trajectory information between key poses, they cannot produce constraint motions such as opening a cabinet. Moreover, they can only predict key poses in the demonstrated order, whereas our approach further produces a set of reusable and re-combinable skills.
@@ III Background
In this section, we summarize foundational work on Gaussian Mixture Models for robotic manipulation. We provide additional details including full equations in Sec. S.1 of the supplementary material.
@@ III-A Gaussian Mixture Model
A Gaussian Mixture Model (GMM) with the Hidden Markov Model (HMM) extension [27] and $K$ Gaussian components is defined as $\mathcal{G}=\{\pi_{k},{\{a_{kl}\}_{l=1}^{K}},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$. $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\Sigma}_{k}$ denote the $k$-th component’s mean and covariance matrix and $\pi_{k}$ denotes its prior probability. $a_{kl}$ is the transition probability from state $k$ to state $l$. In the context of Learning from Demonstrations, GMMs are employed to estimate the joint distribution of a disjoint set of input variables $\mathcal{I}$ and output variables $\mathcal{O}$, from a set of task demonstrations $\mathcal{D}$ as
@@
III-C Task-Parameterization
Task-Parameterized GMMs [6] model the movement from the perspective of multiple coordinate frames, called task-parameters, in one joint GMM [10]. These frames are attached to task-relevant objects but may also include the robot’s frame and a world frame. Let $\boldsymbol{x}(f)$ denote the end-effector position in the frame of task-parameter $f$. We discuss the extension to full poses $\boldsymbol{\xi}(f)$ in Sec. III-D. For the time-driven model and $F$ task-parameters, we model =$\mathcal{I}=\{t\}\text{ and }\mathcal{O}=\{\boldsymbol{x}(f)\}_{f=1}^{F}$. In contrast, for the state-driven model, we define =$\mathcal{I}=\{\boldsymbol{x}(f)\}_{f=1}^{F}\text{ and }\mathcal{O}=\{\dot{\boldsymbol{x}}(f)\}_{f=1}^{F}$. Akin to the vanilla GMM, we fit all parameters jointly using EM. Next, we compute the per-frame marginal model $p_{f}(\mathcal{I},\mathcal{O}\mid\mathcal{G})=\left\{\pi_{k}^{f},\boldsymbol{\mu}_{k}^{f},\boldsymbol{\Sigma}_{k}^{f}\right\}_{k=1}^{K}$ for each frame $1\leq f\leq F$. To adapt to a given set of task parameters $\Theta=\{\boldsymbol{A}_{f},\boldsymbol{b}_{f}\}_{f=1}^{F}$, we transform the per-frame marginals to the world frame as
$\begin{split}p_{f}(\mathcal{I},\mathcal{O}\mid\mathcal{G},\Theta)&=\left\{\pi_{k}^{f},\hat{\boldsymbol{\mu}}^{f}_{k},\hat{\boldsymbol{\Sigma}}^{f}_{k}\right\}_{k=1}^{K},\\ \hat{\boldsymbol{\mu}}^{f}_{k}&=\boldsymbol{A}_{f}\boldsymbol{\mu}_{k}^{f}+\boldsymbol{b}_{f},\\ \hat{\boldsymbol{\Sigma}}^{f}_{k}&=\boldsymbol{A}_{f}\boldsymbol{\Sigma}_{k}^{f}\boldsymbol{A}_{f}^{T}.\end{split}$ (3)
@@ III-C Task-Parameterization
III-D Riemannian Manifolds
We model the robot’s end-effector pose $\boldsymbol{\xi}=\begin{bmatrix}\boldsymbol{x}&\boldsymbol{\theta}\end{bmatrix}^{T}$ on the Riemmanian Manifold $\mathcal{M}_{\text{pose}}=\mathbb{R}^{3}\times\mathcal{S}^{3}$. Whereas GMMs typically assume Euclidean data, the naive Euclidean mean of a set of quaternions is generally not a valid quaternion. Each Riemannian manifold $\mathcal{M}$ is differentiable, i.e., at each point $\mathbf{x}$, it has a tangent space $\mathcal{T}_{\mathbf{x}}\mathcal{M}$, that is a real vector space. Zeestraten [10] formulates all previously Euclidean operations in EM and GMR, including the quaternion center of mass, as non-linear Maximum Likelihood Estimation problems on the tangent space. We provide additional details in Sec. S.2 of the supplementary material and Zeestraten [10] gives an excellent introduction into Riemannian Manifolds. Note that parallel transport is used to implement the frame transformations. We found this to sometimes introduce unwanted correlations into the model’s covariance, leading to faulty predictions. We discuss the problem and our proposed solution in detail in Sec. S.3 of the supplementary material.
@@ IV Technical Approach
Figure 5: Velocity-based skill segmentation on two noisy robot trajectories for a pouring task. Red ellipses indicate candidate segmentation points that were filtered out: the first set is too close to the start of the trajectory before the robot begins moving, while the second set results from noise in the trajectory. The final segmentation points are the centers of extended sub-threshold segments.
@@ IV-A Action Factorization
In Sec. III-B, we discussed that solving manipulation tasks can require modeling the end-effector’s velocity $\dot{\boldsymbol{\xi}}$ and not only its pose $\boldsymbol{\xi}$, e.g. if time-independence is required. Fig. 2 illustrates that naively modeling $\dot{\boldsymbol{\xi}}$ on the manifold $\mathcal{M}_{\text{pose}}=\mathbb{R}^{3}\times\mathcal{S}^{3}$ (introduced in Sec. III-D) is difficult because velocities are not Euclidean. Instead, they are vector quantities defined by a direction and magnitude. We therefore propose to factorize the end-effector velocity $\dot{\boldsymbol{\xi}}$ as
@@
Our insight is three-fold. First, in the task-parameterized setting, only the direction of the velocities varies between different fixed coordinate frames, not their magnitude. Thus, we reduce the dimensionality of the problem by modeling the magnitudes globally. Second, end-effector velocities are not Euclidean due to the curvature of the trajectories and the motion constraints and smooth acceleration of the robot arm. Consequently, the velocity can be modeled more effectively when factorized, as illustrated in Fig. 2. We validate this aspect empirically in Sec. V-B. To model the factorized velocities, we leverage the Riemannian toolbox introduced in Sec. III-D. Third, we leverage the isolated velocities to automatically segment complex task trajectories into the involved skills. We found this to drastically improve the quality and transferability of the learned model and discuss it in detail in Sec. IV-C. We provide additional details on how to implement the frame transformations for $\mathcal{S}^{2}$ in Sec. S.2 of the supplementary material.
@@ IV-C Skill Segmentation
Robot manipulation tasks often involve a set of skills. The pick and place task in Fig. 4 consists of four skills. 1. Align the gripper with the first object. 2. Grasp it. 3. Align the first object with the second object. 4. Place the first object. All skills have distinct dynamics, so it is difficult to model them jointly in a single state-driven GMM. Furthermore, the skills have different relative durations across task instances, leading to poor temporal alignment across demonstrations. Fig. 3 illustrates that, as a consequence, they are difficult to model with a single time-driven GMM.
Therefore, we propose to automatically segment the full demonstration trajectories into a sequence of skill demonstrations and to learn one model per skill. This allows us to temporally align the skills and to learn local dynamics, specific to the skills. The skill model can be time-driven or state-driven. Furthermore, it allows us to reuse the learned skills by sequencing them in novel ways. Lastly, in Sec. IV-E and Sec. IV-F we discuss its synergistic effects with our initialization strategy and our automatic task-parameterization, respectively. To segment the trajectories, we threshold the action magnitudes introduced in Sec. IV-A. We cluster observations with near-zero action magnitudes into time-based segments, yielding a set of candidate segmentation points. We then filter out candidates that are too close to the beginning or end of each trajectory, as well as segments below a minimum length threshold. As shown in Fig. 5, short segments, often only a single time step long, are typically the result of noise in the trajectory, whereas valid segmentation points correspond to extended periods of sub-threshold velocity. We then use the temporal centers of the remaining segments as segmentation points, assuming that skills are presented in the same order across demonstrations. Afterward, we align the skill demonstrations by resampling them to match their average duration. Finally, we fit one TP-HMM per skill.